Two identical charges of 100 microcoulombs each are 12 cm apart. What is the electric potential at a point midway between them? What is the intensity of the electric field at the midway between them?

1
Expert's answer
2022-03-17T09:47:20-0400

Potential

"V=V_1+V_2"

"V=\frac{kq_1}{r}+\frac{kq_2}{r}"


"V=\frac{9\times10^{9}\times100\times10^{-6}}{0.06}+\frac{9\times10^{9}\times100\times10^{-6}}{0.06}=(15+15)\times10^6V"

"V=30\times10^6V"

Electric field

At symmetry

"E=\frac{kq_1}{r^2}-\frac{kq_2}{r^2}"


"E=\frac{9\times10^{9}\times100\times10^{-6}}{0.06^2}-\frac{9\times10^{9}\times100\times10^{-6}}{0.06^2}=(25-25)\times10^7V/m"

"E_{net}=0V\/m"